OpenStudy (love_to_love_you):

Please help! Will fan and medal! A softball thrown into the air has a parabolic trajectory. After reaching a maximum height of 28 feet, the ball covers a ground distance of 4 feet before hitting the ground. What equation describes the ball's trajectory?

OpenStudy (love_to_love_you):

a) x^2 = -7/4y + 49/2 b) x^2 = -4/7y + 16 c) x^2 = -4y + 28 d) x^2 = -1/7y + 112/7

OpenStudy (kikidee123):

im pretty sure its c but i may be wrong

OpenStudy (navyfighter):

You are correct

OpenStudy (love_to_love_you):

I wish I could medal you both. Thank you!

OpenStudy (navyfighter):

You welcome :D

OpenStudy (kikidee123):

no problem